Circular No. 8323

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN OPHIUCHUS H. Yamaoka, Kyushu University, reports the discovery of a possible

nova (mag ∼ 11.1) by Akira Takao (Kitakyushu, Japan) on unfiltered CCD images taken on Apr. 14.80 UT, the position given as α = 17^{h}39^{m}00^{s}, δ = −23^{o}28′.5 (equinox 2000.0; uncertainty ≥ 10′′). Nothing was visible at this location on his images taken on 2002 Mar. 18–2004 Mar. 31 (limiting mag 11.9). S. Nakano, Sumoto, Japan, forwards the following precise position measured by K. Itagaki (Teppo-cho, Japan) from his unfiltered CCD image taken with a 0.60-m reflector on Apr. 15.684 (when the new object appeared at mag 10.2): α = 17^{h}38^{m}45^{s}.49, δ = −23^{o}28′18′′.5. Nothing obvious appears near this position on the Digitized Sky Survey. The closest star in the USNO-A2.0 catalogue is a very red object (red mag 17.2) with position end figures 45^{s}.367, 17′′.56. A. Price, AAVSO, notes that ASAS-3 data show the new object at V = 10.9 on Apr. 14.31.

SUPERNOVA 2001em IN UGC 11794 D. Pooley and W. H. G. Lewin, Massachusetts Institute of Technol-

ogy, report the detection of x-ray emission (154 counts) from the type-Ib/c supernova 2001em (IAUC 7722, 7737, 8282) with the Chandra x-ray observatory. A 14600-s ACIS-S3 observation on Apr. 4.81 UT revealed a source at α = 21^{h}42^{m}23^{s}.60, δ = +12^{o}29′50′′.3 (equinox 2000.0; ± 0′′.5 in each coordinate), in close agreement with the reported optical and radio positions. Preliminary spectral fits (using either a MEKAL or power-law model) indicate a hard spectrum (MEKAL temperature above the model limit of kT = 80 keV or power-law photon index of 1.1) with a fitted column density of N_{H} ≃ 1.6 × 10^{21} cm^{−2}. The unabsorbed 0.5–8-keV luminosity calculated from both models agrees well and is roughly 10^{41} erg/s (for a distance of 80 Mpc). This high x-ray luminosity (the second highest ever reported for a young supernova) could indicate a substantial amount of circumstellar interaction. Further observations at other wavelengths are encouraged.

COMET C/2001 Q4 (NEAT) Recent visual total-magnitude and coma-diameter estimates: Apr. 7.93

UT, 5.3, 14′ (A. Amorim, Florianopolis, Brazil, 7×50 binoculars); 9.45, 5.6, 12′ (A. Pearce, Kalgoorlie, W. Australia, 8×40 binoculars); 10.40, 5.2, – (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars); 13.91, 4.9, 15′ (Amorim); 14.86, 5.0, 11′ (Pearce, Nedlands, W. Australia).

2004 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green